# Exploring Mutual Empowerment Between Wireless Networks and RL-based LLMs: A Survey

### v2

@@ III-D Pure RL-based LLMs
Additionally, DeepSeek-R1 replaces the input prompt with a specific reasoning question format for training the model [28]. This technique allows LLMs to generate a chain of thought (CoT) that can explain the reasoning behind the answer. Prior to providing the answer, LLMs engage in extensive logical deliberation, sometimes even self-correcting with an ”aha moment” [28] before giving the final response. Notably, this technique helps train the reason model without necessitating collected data to train the reward model, thereby enhancing its scalability and adaptability across diverse contexts. However, this approach has its drawbacks, notably in its low generalizability and performance compared to other methods in other tasks [28]. Moreover, the complexity of defining reward rules for specific domains presents a significant obstacle, hindering the seamless application of this approach across various domains.
Figure 6: Motivations, challenges, and potential solutions for empowering wireless communication systems with LLMs.
Figure 7: Challenges for empowering wireless communication systems with LLMs.
IV RL-based LLMs for Wireless Networks
@@
IV-A Motivations for Empowering Wireless Networks with RL-based LLMs
@@
Strong Generalization of LLMs. Wireless networks operate in highly dynamic and complex environments where network conditions, user demands, and interference patterns fluctuate over time [51, 142]. For instance, network congestion can vary depending on the number of users in a specific area, while interference may be influenced by geographic location and environmental factors [143, 144, 145]. Similarly, user demands, such as data rates or latency requirements, can change throughout the day [146]. In these situations, traditional machine learning models often struggle to adapt, requiring frequent retraining with newly collected data, which can be costly and inefficient.
@@
In contrast, LLMs possess strong generalization capabilities, enabling them to effectively manage diverse and evolving situations without the need for constant retraining [147]. Generalization refers to the ability to make effective decisions and predictions in unseen domains without requiring retraining [148]. This capability is enabled by leveraging knowledge from large-scale pretraining, which allows LLMs to transfer insights across different domains, thereby enhancing both efficiency and adaptability. For instance, an LLM trained on a wide range of textual data can be easily adapted to tasks such as medical diagnosis [149], sentiment analysis for social media [150], or machine translation [151] with minimal fine-tuning.
Therefore, this generalization ability is especially valuable in wireless networks, where conditions are constantly evolving, and the ability to quickly adapt to new scenarios is crucial for maintaining efficient and reliable performance.
@@
Shortage of High-Quality and Structured Data in Wireless Networks. Wireless networks often face the challenge of obtaining high-quality, structured data due to the dynamic nature of the environment and the decentralized distribution of data sources, particularly in the era of big data [152, 153]. For example, wireless communication systems rely on a wide variety of sensors, such as base stations, user equipment, and environmental factors like weather conditions, which produce vast amounts of unstructured and noisy data [154, 155, 156]. This data is often incomplete, inconsistent, and challenging to organize into a usable form for analysis or model training. Moreover, the vast amount of data generated across different devices and locations makes it difficult to collect and label the necessary datasets for traditional machine learning approaches.
@@
LLMs, on the other hand, can potentially address this issue by leveraging their ability to process and make sense of unstructured data [157]. By pretraining on large-scale data, LLMs can extract meaningful patterns and relationships from vast amounts of raw, unstructured data without the need for explicit labeling or structured data [158]. For instance, LLMs can be trained on logs from wireless networks or communication traces, where they learn to identify patterns of network congestion, interference, and traffic fluctuations, even when the data is noisy or incomplete [159]. With minimal fine-tuning, LLMs can apply these insights to improve network resource allocation, optimize routing protocols, or predict future demand without requiring the extensive amount of structured, labeled data traditionally needed for such tasks. The proof-of-concept results from WirelessLLM [81] highlight the effectiveness of LLMs in understanding and addressing real-world wireless communication scenarios, such as network monitoring, resource management, and protocol understanding.
@@
This ability to work with noisy, unstructured data and still deliver reliable predictions makes LLMs a promising solution for the data scarcity challenges inherent in wireless networks. Moreover, LLMs can also generate synthetic data to supplement the scarcity of high-quality, structured data in wireless networks [160, 161]. By leveraging their generative capabilities, LLMs can simulate realistic network scenarios, creating data that reflects various conditions, such as different levels of interference, congestion, or traffic patterns. This synthetic data can be used to train models, test algorithms, or augment existing datasets, thereby overcoming data limitations and improving the robustness of wireless network systems.
@@
Intelligent Decision-Making Requirements in Wireless Networks. Wireless networks are subject to dynamic and unpredictable conditions, such as interference, congestion, and mobility of devices [162, 163]. These constantly changing factors make it difficult to predict traffic patterns and optimize resources effectively. Meanwhile, effective resource management, such as spectrum allocation, power control, and user scheduling, is essential to ensure optimal network performance. However, these tasks require continuous real-time adaptation and decision-making, as network conditions change rapidly and are often difficult to predict. Traditional methods, such as heuristic-based user scheduling [164], rule-based scheduling [165], and static power control algorithms [166], often fail to adapt to the dynamic nature of the network [1].
@@
To overcome the limitations of traditional methods, which often suffer from unreliable decision-making, LLMs present a compelling alternative by leveraging their inference capabilities to process complex, unstructured data and make intelligent, context-aware decisions [167]. For instance, an LLM can analyze both historical trends and real-time network data to predict congestion and proactively adjust resource allocation, such as reallocating spectrum or modifying transmission power to prevent network overload. [80] demonstrates that leveraging the inference capabilities of LLMs can achieve performance comparable to traditional reinforcement learning while bypassing extensive model training and hyperparameter tuning. Moreover, when integrated with RL, LLMs enable dynamic optimization of resource management by incorporating historical data, real-time inputs, and human feedback to adapt to evolving network conditions. By continuously learning from human input and channel feedback, the system can update itself in real-time, ensuring more responsive and adaptive network management.
@@
Adaptable and Reasoning in LLMs. Since LLMs are trained on diverse data sources, they demonstrate remarkable adaptability, enabling them to be customized for specific domains through techniques such as fine-tuning [168], transfer learning [169, 170], prompt engineering [171], and retrieval-augmented generation (RAG) [172]. These techniques enable LLMs to quickly become domain-specific experts, understand complex tasks, and respond effectively to specialized queries. Specifically, fine-tuning allows LLMs to optimize predictions based on domain-specific nuances in downstream tasks, making them highly effective in those contexts. Transfer learning enables LLMs to transfer knowledge from one domain or task to another, leveraging patterns and features learned from the source domain to adapt more rapidly to new tasks or domains with less data. Similarly, prompt engineering allows users to design tailored instructions that guide the behavior of the LLM, ensuring it produces relevant and precise responses within the target domain. Finally, the integration of RAG further enhances adaptability by enabling access to external knowledge sources, ensuring up-to-date and context-aware decision-making. This adaptability empowers the model to continuously adjust and optimize its performance across domains, minimizing the need for extensive retraining.
Figure 8: An example of CoT. Different from standard prompting, CoT prompting explicitly outlines the reasoning process, leading to more interpretable and accurate results.
@@
In addition to adaptability, the latest LLMs now possess advanced reasoning capabilities, such as chain of thought (CoT) [173] methods found in models like DeepSeek-R1 [28] and GPT-4o [135]. The CoT method provides transparent and logical thought processes, enabling the models to break down complex problems into simpler components and reason through intricate scenarios step by step. For instance, CoT can explain why the sea appears blue by detailing how water molecules preferentially absorb longer wavelengths (e.g., red) while scattering shorter wavelengths (e.g., blue), resulting in the characteristic blue color of the ocean. Fig. 8 illustrates an example of the reasoning process. By elucidating their reasoning chain, LLMs not only demystify challenging problems but also offer valuable insights for decision-making. This capacity to deconstruct complex tasks into understandable steps is particularly critical in managing the uncertainty and variability inherent in wireless networks. By simplifying complex tasks and making intelligent predictions without the need for extensive retraining, LLMs can optimize resource allocation and enhance network management, thereby providing effective solutions to the dynamic challenges of wireless environments.
IV-B Challenges
@@
High Computing Resource Requirements. LLMs are structured on transformer architecture, known for their quadratic complexity. Given that wireless networks necessitate ultra-low latency for communication, striking a balance between accuracy and complexity becomes imperative. Researchers and developers are tasked with crafting efficient models that can navigate this challenge effectively. Moreover, the consideration of power consumption is a necessity, particularly for devices such as IoT or mobile devices. This additional constraint further complicates the optimization of models for real-world deployment within resource-constrained environments.
@@
Poor Interpretability of LLMs. Although there are efforts to develop tools and techniques for explaining and interpreting LLMs, these resources are still in their early stages of development. Existing methods, such as attention mechanisms and feature visualization, provide some insights but are often limited in fully explaining the model’s behavior. Consequently, it is important to develop techniques that can explain how LLMs generate outputs, facilitating the regulation of these models’ behavior for industrial applications. Moreover, these techniques can empower researchers to enhance their understanding of LLMs, enabling them to tackle limitations and boost performance effectively.
Figure 9: Potential solutions for empowering wireless communication systems with LLMs.
@@
Instability Induced by Dynamic Changes in Wireless Networks. The interactions among various parameters and layers in LLMs are non-linear, meaning that small changes in input can result in significant and unpredictable output changes. This makes it difficult to predict and interpret the model’s behavior across diverse scenarios. The vast number of parameters and layers involved within these models poses a challenge in tracing and understanding their decision-making process. In wireless communication, this complexity is further compounded by the dynamic and heterogeneous nature of network environments and the need to process and analyze vast amounts of unstructured data available on most network devices. Additionally, the application of semantic communication, which aims to understand and interpret the meaning behind the transmitted data rather than just the raw bits, introduces further challenges. When only the context is conveyed in the message, the inputs to LLMs may vary, yet they must produce consistent outputs based on this shared context.
@@
Data Privacy and Security Issues. LLMs require large datasets for training to build a robust model for wireless networks. However, these data samples may contain user-sensitive information that can not be leaked outside the network. It is imperative to anonymize and secure this data during collection and distribution to forbid unauthorized access and potential abuse. Furthermore, LLMs can be vulnerable to exploitation, potentially exposing user information and offering adversaries fresh avenues for attack. Therefore, guaranteeing the accuracy and integrity of the data employed by LLMs is significant in maintaining communication security.
IV-C Potential Solutions

### v3

@@ 2.2.3. Pure RL-based LLMs
Another innovative approach involves replacing the reward model with predefined reward rules, as proposed by DeepSeek-R1 (Guo et al., 2025), which has demonstrated exceptional performance in reasoning tasks for LLMs. This method eliminates the need for supervised fine-tuning, relying solely on RL to guide model optimization. Notably, recent findings from DeepSeek-R1-Zero provide the first evidence that reasoning behaviors can naturally emerge through pure RL (Guo et al., 2025), highlighting its potential as a powerful alternative to conventional training paradigms. Specifically, instead of relying on a pre-trained reward model derived from a rating dataset, DeepSeek-R1 implements a rule-based reward system comprising two primary types of rewards: accuracy rewards and format rewards. Accuracy rewards are designed to enable the model to assess the correctness of its responses. For instance, in scenarios such as math problems with deterministic solutions, the model can only earn a point if it provides the correct answer. On the other hand, format rewards force the model to respond in accordance with a predefined format to be considered a valid response and earn a point. In addition, DeepSeek-R1 replaces the input prompt with a specific reasoning question format for training the model (Guo et al., 2025). This technique allows LLMs to generate a chain of thought (CoT) that can explain the reasoning behind the answer. Fig. 5 illustrates an example of the reasoning process. Prior to providing the answer, LLMs engage in extensive logical deliberation, sometimes even self-correcting with an "aha moment" (Guo et al., 2025) before giving the final response. Notably, this technique helps train the reason model without necessitating collected data to train the reward model, thereby enhancing its scalability and adaptability across diverse contexts. However, this approach also has its drawbacks, notably in its low generalizability and performance compared to other methods in other tasks (Guo et al., 2025). Moreover, the complexity of defining reward rules for specific domains presents a significant obstacle, hindering the seamless application of this approach across various domains. As a result, while pure RL approaches offer a promising and scalable alternative to traditional training pipelines, further research is required to improve their adaptability, robustness, and applicability to a wider range of real-world scenarios (Xu et al., 2025a; Guo et al., 2025).
2.3. Wireless Networks Meet LLMs
Although significant progress has been made in optimizing wireless networks through advanced resource management and intelligent architectures, these solutions often rely on task-specific algorithms that lack generalization and adaptability (Chen et al., 2024d). In contrast, the emergence of LLMs, such as DeepSeek (Guo et al., 2025), ChatGPT (Achiam et al., 2023), and Gemini (Team et al., 2023), has introduced a powerful new paradigm for cross-domain intelligence and decision-making. With their exceptional capabilities in understanding, reasoning, and multimodal interaction, LLMs have the potential to fundamentally transform how wireless networks are designed, managed, and optimized. In particular, LLMs enable learning from heterogeneous data sources, extracting insights from complex network states, and controlling network elements through multimodal inputs. Conversely, wireless communication infrastructures play a critical role in supporting the efficient development and scalable deployment of LLMs. This bidirectional relationship opens up promising opportunities for building LLM-driven wireless frameworks, while simultaneously leveraging wireless advancements to enable more intelligent and efficient LLM systems. Despite these emerging possibilities, there remains a lack of contemporary surveys exploring the mutual empowerment of these two rapidly evolving fields. Therefore, in the following section, we examine the synergy between wireless networks and LLMs from three perspectives: motivations, challenges, and potential solutions. This discussion lays the foundation for a deeper understanding of how the integration of LLMs can unlock new levels of intelligence and automation in future wireless systems, and conversely, how advances in wireless networks can enhance the capabilities and deployment of LLMs. It is worth noting that since most popular LLMs (e.g., ChatGPT, Gemini, Grok, and DeepSeek) incorporate RL in their training process, we use the terms RL-based LLMs and LLMs interchangeably throughout this paper.
3. RL-based LLMs for Wireless Networks
@@
3.1. Motivations for Empowering Wireless Networks with LLMs
3.1.1. Strong Generalization of LLMs
@@
Wireless networks operate in highly dynamic and complex environments where network conditions, user demands, and interference patterns fluctuate over time (Letaief and Zhang, 2006; Mao et al., 2018). For instance, network congestion can vary depending on the number of users in a specific area, while interference may be influenced by geographic location and environmental factors (Adhikary et al., 2023c, 2024b; Wei et al., 2020). Similarly, user demands, such as data rates or latency requirements, can change throughout the day (Jiang et al., 2018). In these situations, traditional machine learning models often struggle to adapt, requiring frequent retraining with newly collected data, which can be costly and inefficient.
@@
In contrast, LLMs possess strong generalization capabilities, enabling them to effectively manage diverse and evolving situations without the need for constant retraining (Zheng et al., 2024). Generalization refers to the ability to make effective decisions and predictions in unseen domains without requiring retraining (Raha et al., 2024). This capability is enabled by leveraging knowledge from large-scale pretraining, which allows LLMs to transfer insights across different domains, thereby enhancing both efficiency and adaptability. For instance, an LLM trained on a wide range of textual data can be easily adapted to tasks such as medical diagnosis (Shi et al., 2024), sentiment analysis for social media (Kumar et al., 2023), or machine translation (Eschbach-Dymanus et al., 2024) with minimal fine-tuning. Therefore, this generalization ability is especially valuable in wireless networks, where conditions are constantly evolving, and the ability to adapt to new scenarios quickly is crucial for maintaining efficient and reliable performance.
Figure 6. Motivations, challenges, and potential solutions for empowering wireless communication systems with LLMs.
3.1.2. Shortage of High-Quality and Structured Data in Wireless Networks
@@
Wireless networks often face the challenge of obtaining high-quality, structured data due to the dynamic nature of the environment and the decentralized distribution of data sources, particularly in the era of big data (Sharma and Wang, 2017; Nguyen et al., 2020). For example, wireless communication systems rely on a wide variety of sensors, such as base stations, user equipment, and environmental factors like weather conditions, which produce vast amounts of unstructured and noisy data (Dai et al., 2019; Nguyen et al., 2022; Garg, 2010). This data is often incomplete, inconsistent, and challenging to organize into a usable form for analysis or model training. Moreover, the vast amount of data generated across different devices and locations makes it difficult to collect and label the necessary datasets for traditional machine learning approaches.
@@
LLMs, on the other hand, can potentially address this issue by leveraging their ability to process and make sense of unstructured data (Vijayan, 2023). By pretraining on large-scale data, LLMs can extract meaningful patterns and relationships from vast amounts of raw, unstructured data without the need for explicit labeling or structured data (Dagdelen et al., 2024). For instance, LLMs can be trained on logs from wireless networks or communication traces, where they learn to identify patterns of network congestion, interference, and traffic fluctuations, even when the data is noisy or incomplete (Kan et al., 2024). With minimal fine-tuning, LLMs can apply these insights to improve network resource allocation, optimize routing protocols, or predict future demand without requiring the extensive amount of structured, labeled data traditionally needed for such tasks. The proof-of-concept results from WirelessLLM (Shao et al., 2024b) highlight the effectiveness of LLMs in understanding and addressing real-world wireless communication scenarios, such as network monitoring, resource management, and protocol understanding.
@@
This ability to work with noisy, unstructured data and still deliver reliable predictions makes LLMs a promising solution for the data scarcity challenges inherent in wireless networks. Moreover, LLMs can also generate synthetic data to supplement the scarcity of high-quality, structured data in wireless networks (Long et al., 2024; Tao et al., 2024). By leveraging their generative capabilities, LLMs can simulate realistic network scenarios, creating data that reflects various conditions, such as different levels of interference, congestion, or traffic patterns. This synthetic data can be used to train models, test algorithms, or augment existing datasets, thereby overcoming data limitations and improving the robustness of wireless network systems.
3.1.3. Intelligent Decision-Making Requirements in Wireless Networks
@@
Wireless networks are subject to dynamic and unpredictable conditions, such as interference, congestion, and mobility of devices (Mobile, 2002; Praveen and Kumanan, 2024). These constantly changing factors make it difficult to predict traffic patterns and optimize resources effectively. Meanwhile, effective resource management, such as spectrum allocation, power control, and user scheduling, is essential to ensure optimal network performance. However, these tasks require continuous real-time adaptation and decision-making, as network conditions change rapidly and are often difficult to predict. Traditional methods, such as heuristic-based user scheduling (Zhong et al., 2019), rule-based scheduling (Selvi et al., 2019), and static power control algorithms (Pantazis and Vergados, 2007), often fail to adapt to the dynamic nature of the network (Liu et al., 2024a).
@@
To overcome the limitations of traditional methods, which often suffer from unreliable decision-making, LLMs present a compelling alternative by leveraging their inference capabilities to process complex, unstructured data and make intelligent, context-aware decisions (Qiu et al., 2024b). For instance, an LLM can analyze both historical trends and real-time network data to predict congestion and proactively adjust resource allocation, such as reallocating spectrum or modifying transmission power to prevent network overload. (Zhou et al., 2024a) demonstrates that leveraging the inference capabilities of LLMs can achieve performance comparable to traditional reinforcement learning while bypassing extensive model training and hyperparameter tuning. Moreover, when integrated with RL, LLMs enable dynamic optimization of resource management by incorporating historical data, real-time inputs, and human feedback to adapt to evolving network conditions. By continuously learning from human input and channel feedback, the system can update itself in real time, ensuring more responsive and adaptive network management.
3.1.4. Adaptable and Reasoning in LLMs
@@
Since LLMs are trained on diverse data sources, they demonstrate remarkable adaptability, enabling them to be customized for specific domains through techniques such as fine-tuning (Xin et al., 2024), transfer learning (Adhikary et al., 2024a; Tan et al., 2018), prompt engineering (Sahoo et al., 2024), and retrieval-augmented generation (RAG) (Yu et al., 2024). These techniques enable LLMs to quickly become domain-specific experts, understand complex tasks, and respond effectively to specialized queries. Specifically, fine-tuning allows LLMs to optimize predictions based on domain-specific nuances in downstream tasks, making them highly effective in those contexts. Transfer learning enables LLMs to transfer knowledge from one domain or task to another, leveraging patterns and features learned from the source domain to adapt more rapidly to new tasks or domains with less data. Similarly, prompt engineering allows users to design tailored instructions that guide the behavior of the LLM, ensuring it produces relevant and precise responses within the target domain. Finally, the integration of RAG further enhances adaptability by enabling access to external knowledge sources, ensuring up-to-date and context-aware decision-making. This adaptability empowers the model to continuously adjust and optimize its performance across domains, minimizing the need for extensive retraining.
@@
In addition to adaptability, the latest LLMs now possess advanced reasoning capabilities, such as chain of thought (CoT) (Wei et al., 2022b) methods found in models like DeepSeek-R1 (Guo et al., 2025) and GPT-4o (Hurst et al., 2024). The CoT method provides transparent and logical thought processes, enabling the models to break down complex problems into simpler components and reason through intricate scenarios step by step. For instance, CoT can explain why the sea appears blue by detailing how water molecules preferentially absorb longer wavelengths (e.g., red) while scattering shorter wavelengths (e.g., blue), resulting in the characteristic blue color of the ocean. By elucidating their reasoning chain, LLMs not only demystify challenging problems but also offer valuable insights for decision-making. This capacity to deconstruct complex tasks into understandable steps is particularly critical in managing the uncertainty and variability inherent in wireless networks. By simplifying complex tasks and making intelligent predictions without the need for extensive retraining, LLMs can optimize resource allocation and enhance network management, thereby providing effective solutions to the dynamic challenges of wireless environments.
Figure 7. Challenges for empowering wireless communication systems with LLMs.
3.2. Challenges
@@ 3.2.1. High Computing Resource Requirements
LLMs are structured on transformer architecture, known for their quadratic complexity (Keles et al., 2023). Given that wireless networks necessitate ultra-low latency for communication, striking a balance between accuracy and complexity becomes imperative (Sutton et al., 2019). Researchers and developers are tasked with crafting efficient models that can navigate this challenge effectively. Moreover, the consideration of power consumption is a necessity, particularly for devices such as IoT or mobile devices. This additional constraint further complicates the optimization of models for real-world deployment within resource-constrained environments.
3.2.2. Poor Interpretability of LLMs
@@
Although there are efforts to develop tools and techniques for explaining and interpreting LLMs (Yang et al., 2023b), these resources are still in their early stages of development. Existing methods, such as AttentionViz (Yeh et al., 2024) and feature visualization (Wang et al., 2021), provide some insights but are often limited in fully explaining the model’s behavior. Consequently, it is important to develop techniques that can explain how LLMs generate outputs, facilitating the regulation of these models’ behavior for industrial applications. Moreover, these techniques can empower researchers to enhance their understanding of LLMs, enabling them to tackle limitations and boost performance effectively.
3.2.3. Instability Induced by Dynamic Changes in Wireless Networks
@@
The interactions among various parameters and layers in LLMs are non-linear, meaning that small changes in input can result in significant and unpredictable output changes (Kumar, 2024). This makes it difficult to predict and interpret the model’s behavior across diverse scenarios. The vast number of parameters and layers (Guo et al., 2025; Hurst et al., 2024; Reyes-Rivera et al., 2024) involved within these models poses a challenge in tracing and understanding their decision-making process. In wireless communication, this complexity is further compounded by the dynamic and heterogeneous nature of network environments and the need to process and analyze vast amounts of unstructured data available on most network devices (Xu et al., 2021). Additionally, the application of semantic communication (Nguyen et al., 2024b), which aims to understand and interpret the meaning behind the transmitted data rather than just the raw bits, introduces further challenges. When only the context is conveyed in the message, the inputs to LLMs may vary, yet they must produce consistent outputs based on this shared context.
3.2.4. Data Privacy and Security Issues
@@
LLMs require large datasets for training to build a robust model for wireless networks (Guo et al., 2025; Hurst et al., 2024; Reyes-Rivera et al., 2024). However, these data samples may contain user-sensitive information that can not be leaked outside the network (Yan et al., 2024). It is imperative to anonymize and secure this data during collection and distribution to forbid unauthorized access and potential abuse. Furthermore, LLMs can be vulnerable to exploitation, potentially exposing user information and offering adversaries fresh avenues for attack (Das et al., 2025). Therefore, guaranteeing the accuracy and integrity of the data employed by LLMs is significant in maintaining communication security.
3.3. Potential Solutions
3.3.1. Enabling Edge Computing for LLM Training and Inference